# Local logistic regression for dimension reduction in classification

### v1

@@ 2.2 Nearest-neighbor penalized local logistic estimator
These results provide statistical guarantees on the estimation of directions belonging to the dimension reduction subspace, and will therefore be useful in order to find such guarantees on the dimension reduction procedure. The rate obtained in Corollary 3 has also been obtained as part of non-asymptotic bounds on nearest neighbor estimation of the gradient of the regression function, although this of course requires strong assumptions on the tail behavior of the noise, see e.g. Theorem 1 in Ausset et al. (2021). The best achievable rate of convergence, obtained for $k=n^{4/(p+4)}$, is $n^{-1/(p+4)}$ for the estimation of the gradient vector. This rate is optimal, see Stone (1982), is the same as the one obtained in Ausset et al. (2021), and is better than the rate $n^{-1/(3(p+4))}$ obtained in Kang and Shin (2022).
The above results only require $X$ to have a continuous density that is positive around $x$, and the conditional probability to be twice continuously differentiable at $x$. As a consequence the best that can be hoped for is, of course, a pointwise asymptotic convergence result in the spirit of Theorem 2. Stronger results, such as uniform convergence results for $\widehat{b}_{n}(x)$ over compact subsets of the support of $X$, could be obtained under much more restrictive conditions.
@@ 2.3 Aggregating the directions
The optimization procedure (1) allows to estimate $\nabla\ell(x)$ which, as explained in Section 2.1, belongs to the central subspace for each $x\in S_{X}$. This subspace can be recovered by estimating several such directions at different points $x$. To this aim, given a probability measure $\mu$ supported on $S_{X}$, let
@@
Suppose that $\mu$ is a finitely supported measure on $S_{X}$ and that for each $x\in\operatorname{supp}(\mu)$, assumptions (A1) and (A2) are satisfied. If $k:=k_{n}\to\infty$ and $\lambda:=\lambda_{n}$ are such that $k^{1+p/2}/n\to\infty$, $k^{1+p/4}/n$ is bounded and $n\lambda^{p}/k^{1+p/2}$ converges to a finite constant, we have
@@
The above rate of convergence can be extended to the eigenprojector $\widehat{P}_{\widehat{\beta}}$ as indicated by Lemma 4.1 in Tyler (1981). Also note that a natural choice for $\mu$, which is also the one made in our numerical experiments, is the empirical measure of $X_{1},\ldots,X_{n}$. This is, of course, a random measure whose number of atoms is not bounded with respect to $n$. From a theoretical perspective, we conjecture, following results given in Hristache et al. (2001) and Dalalyan et al. (2008), that such a choice would lead to a different rate of convergence compared to the one given in the above corollary, under substantially stronger regularity assumptions than ours.
@@ 3.1.1 Estimation of the dimension reduction matrix
For a given choice of $m$ (the number of $b$’s used to estimate $M$), a given value of $\lambda$, the penalization parameter in the optimization problem, and $k$ the number of neighbors, the computation of the dimension reduction matrix is straightforward, see Algorithm 1 below (the optimization is carried out using the R function glmnet from the package of the same name).
@@
To estimate the directions featured in (2), we use either the pure nearest-neighbor logistic log-likelihood without penalization or its penalized version. For the latter, we require an optimal choice of $\lambda$ which balances the regularization and goodness-of-fit. We propose to select the same parameter $\lambda$ for all $x\in\mathcal{X}_{m}$ in order to alleviate the computational burden. We therefore select $\lambda$ by 10-fold cross-validation at the average point $x=\overline{x}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$. In other words, we divide the data into 10 randomly selected subsets of equal size, each subset serving as a validation set while the remaining subsets are used for model fitting. The fitted model is assessed on the validation set using the misclassification error, that is, the proportion of observations whose label is not correctly predicted, which is nothing but the empirical counterpart of the misclassification risk $\mathcal{R}(g)=\mathbb{P}(g(X)\neq Y)$ for a given classifier $g$; in this cross-validation procedure, an observation is labeled as 1 if and only if its predicted probability of success by the nearest-neighbor logistic log-likelihood estimator $\widehat{\pi}_{n}(\overline{x})=\operatorname{expit}(\widehat{a}_{n}(\overline{x}))$ exceeds $1/2$. This evaluation of the quality of the fitted model is then done across a sequence of $\lambda$ values: more precisely, we use the R function cv.glmnet from the glmnet package with type.measure=class, which automatically computes the average misclassification error across all validation sets, and the regularization parameter $\lambda$ selected is the one minimizing this error.
@@ 3.1.2 Selecting the dimension of the reduction subspace
Unlike some dimension reduction approaches such as that of the seminal paper by Li (1991), we estimate the dimension of the reduction subspace using cross-validation on the underlying prediction problem, and not from a test on the eigenvalues of the outer product of gradients $\widehat{M}$. We refer to Bura and Yang (2011) and Portier and Delyon (2014) for more details on the rank testing procedures used to estimate the dimension of the reduction subspace. Our procedure first divides the data into a training set $(X_{\mathrm{train}},Y_{\mathrm{train}})$ and a testing set $(X_{\mathrm{test}},Y_{\mathrm{test}})$. The matrix $\widehat{M}$ and its $p$ orthogonal eigenvectors $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{p}$, in decreasing order according to their eigenvalues, are estimated from the training set following the procedure described in Algorithm 1. For every $d\in\{1,\ldots,p\}$, the first $d$ eigenvectors are then gathered in a matrix $\widehat{\beta}_{(1:d)}\in\mathbb{R}^{p\times d}$, the sets of covariates $X_{\mathrm{train}}$ and $X_{\mathrm{test}}$ are projected onto the sets of lower-dimensional covariates $X_{\mathrm{train}}\widehat{\beta}_{(1:d)}$ and $X_{\mathrm{test}}\widehat{\beta}_{(1:d)}$, and a classifier is learned based on the training set $(X_{\mathrm{train}}\widehat{\beta}_{(1:d)},Y_{\mathrm{train}})$. We use here mainly the knn nearest-neighbor classifier, that is, at a given point $x$, the result of the majority vote among the nearest neighbors of $x$ within the space of projections of the covariates in the training set (with ties broken at random). For this classifier, the training step merely consists of storing the covariates in the training set along with their labels, which will form the basis for the vote at each point. In the real data analysis, we shall also compare our results with the Random Forest classifier, whose training step is nontrivial. The performance of the chosen classifier is then evaluated on the test set $(X_{\mathrm{test}}\widehat{\beta}_{(1:d)},Y_{\mathrm{test}})$, and the dimension retained is the one for which the classifier has the lowest misclassification risk. This forward iterative procedure, summarized in Algorithm 2, makes it possible to strike a balance between dimension reduction and the preservation of relevant information within the dataset. We shall compare the results obtained with the situation where the correct dimension of the reduction subspace is known in order to assess the influence of the dimension selection step.
@@ 3.2 Simulation study
In all the examples presented below, $(X_{i},Y_{i})_{i=1,\ldots,n}$ is a collection of independent and identically distributed random copies of the pair $(X,Y)$, and $X$ is a vector of independent centered and unit Gaussian covariates with dimension $p\geq 8$. We consider the following four examples:
Example 1. The binary response $Y$ has a Bernoulli distribution with parameter $\operatorname{expit}(X_{1})$.
Example 2. The binary response is $Y=\operatorname{sign}\{\sin(X_{1})+X_{2}^{2}+0.2\epsilon\}$, where $\epsilon\sim\mathcal{N}(0,1)$.
Example 3. The binary response is $Y=\operatorname{sign}\{(X_{1}+0.5)(X_{2}-0.5)^{2}+0.2\epsilon\}$, where $\epsilon\sim\mathcal{N}(0,1)$.
Example 4. The binary response is $Y=\operatorname{sign}\{\log(X_{1}^{2})(X_{2}^{2}+X_{3})+0.2\epsilon\}$, where $\epsilon\sim\mathcal{N}(0,1)$.
Example 1 is the simplest instance of a logistic regression model with a single relevant feature, while Examples 2, 3, and 4 are closely related to examples considered by Meng et al. (2020). We compare the proposed approach, with penalization (denoted by LLO$(\lambda>0)$) and without penalization (called LLO$(\lambda=0)$) with three existing competitors: SAVE and PHD, which are inverse regression techniques (Cook and Weisberg, 1991; Li, 1992) (both using the dr function from the R package of the same name) and POTD, relying on optimal transport (Meng et al., 2020). These methods are compared using, on the one hand, the distance between the estimated central subspace and the true central subspace using the Frobenius distance between the projection matrices on these subspaces, that is, $d(\mathcal{S}(\widehat{\beta}),\mathcal{S}({\beta}))=\|P_{\widehat{\beta}}-P_{\beta}\|_{F}$ with $P_{\beta}=\beta\beta^{T}$ for any orthogonal matrix $\beta$ of $p-$dimensional vectors. In each example, the central subspace is explicit: in Example 1, the central subspace is spanned by the first vector of the canonical basis in $\mathbb{R}^{p}$, it is spanned by the first two vectors of this basis in Examples 2 and 3, and it is spanned by the first three vectors of this basis in Example 4. As such, Examples 1 to 4 are ranked in order of complexity of the dimension reduction problem.

### v2

@@ 2.2 Nearest-neighbor penalized local logistic estimator
These results will be useful in order to find statistical guarantees on the dimension reduction procedure. The best achievable rate of convergence, obtained for $k=n^{4/(p+4)}$, is $n^{-1/(p+4)}$ for the estimation of the gradient vector. This rate is optimal, see Stone (1982), and is better than the rate $n^{-1/(6(p+4))}$ obtained in Kang and Shin (2022). A similar (non-asymptotic) bound is obtained in Ausset et al. (2021) in the simpler case of a local least squares estimator of the gradient which is not appropriate in the classification problem.
The above results only require regularity conditions at the point $x$. As a consequence, the best that can be hoped for is a pointwise asymptotic convergence result in the spirit of Theorem 1. Stronger results, such as uniform convergence results for $\widehat{b}_{n}(x)$ over compact subsets of the support of $X$, could be obtained under much more restrictive conditions.
Let us point out that from the estimator $(\widehat{a}_{n}(x),\widehat{b}_{n}(x))$, one can easily construct, using the standard plug-in rule, an estimator of $(\pi(x),\nabla\pi(x))$. This estimator will inherit the good statistical properties of $(\widehat{a}_{n}(x),\widehat{b}_{n}(x))$ established before.
@@ 2.3 Aggregating the directions
The optimization procedure (1) allows to estimate $\nabla\ell(x)$, which belongs to the central subspace for each $x\in S_{X}$. This subspace can be recovered by estimating such directions at different points $x$. To this aim, given a probability measure $\mu$ supported on $S_{X}$, let
@@
Suppose that $\mu$ is a finitely supported measure on $S_{X}$ and that for each $x\in\operatorname{supp}(\mu)$, assumptions (A1) and (A2) are satisfied. If $k:=k_{n}\to\infty$ and $\lambda:=\lambda_{n}$ are such that $k^{1+p/2}/n\to\infty$, $k^{1+p/4}/n$ is bounded and $n\lambda^{p}/k^{1+p/2}$ converges to a finite constant, then
@@
The eigenprojector $\widehat{P}_{\widehat{\beta}}$ has the same rate of convergence, see Lemma 4.1 in Tyler (1981). Also note that a natural choice for $\mu$, which is also the one made in our numerical experiments, is the empirical measure of $X_{1},\ldots,X_{n}$. This is, of course, a random measure whose number of atoms is not bounded with respect to $n$. From a theoretical perspective, we conjecture, following results given in Hristache et al. (2001) and Dalalyan et al. (2008), that such a choice would lead to a different rate of convergence compared to the one given in the above corollary, under substantially stronger regularity assumptions than ours.
2.4 Extension to multi-class generalized linear models
Our general asymptotic theory also applies to much more general settings where $Y$ represents a class label chosen among a finite number of classes, and in generalized linear model (GLM)-type specifications using link functions that may not be equal to the simple logit link. We outline here how this may be done under our framework but omit the technical details for the sake of brevity.
Suppose for the moment that $Y\in E$ is a response variable living in the finite set $E=\{0,1,\ldots,N\}$, with $N\geq 1$, and that the conditional distribution of $Y$ given $X=x$ is part of the one-parameter exponential family, say with conditional probability mass function $y\mapsto h(y)\exp(y\theta(x)-C(\theta(x)))$, where $\theta$ is a univariate parameter mapping and $C$ is a known smooth, strictly monotonic and convex function. Then it is known that the conditional mean $m(x)=\mathbb{E}(Y|X=x)$ satisfies $m(x)=C^{\prime}(\theta(x))$. A GLM-type specification would then impose that $m(x)=\mathbb{E}(Y|X=x)=g(\beta^{T}x)$, for some unknown $\beta\in\mathbb{R}^{p\times d}$, $d\leq p$ and a measurable function $g:\mathbb{R}^{d}\to[0,N]$, so that $\theta(x)=C^{\prime-1}\circ g(\beta^{T}x)$ is a function of $\beta^{T}x$ only. When $g=g_{0}$ is known, the corresponding maximum log-likelihood estimator of $\beta$ is
$\operatorname*{arg\,max}_{b\in\mathbb{R}^{p\times d}}\sum_{i=1}^{n}Y_{i}\Phi_{0}(b^{T}X_{i})-C(\Phi_{0}(b^{T}X_{i})),\mbox{ where }\Phi_{0}=C^{\prime-1}\circ g_{0}.$
Arguing in the same way as in Section 2.2 and viewing the quantity $\Phi(\beta^{T}X)=C^{\prime-1}\circ g(\beta^{T}X)$ as the target of a local linear estimation procedure when $g$ is considered unknown, one may define a nearest-neighbor, penalized local linear GLM-type estimator as
$(\widehat{a}_{n}(x),\widehat{b}_{n}(x))=\operatorname*{arg\,max}_{(a,b)\in\mathbb{R}\times\mathbb{R}^{p}}\{\mathcal{L}_{n}(a,b)-\lambda\|b\|_{1}\}$
with
$\mathcal{L}_{n}(a,b)=\sum_{i\in N_{k}(x)}Y_{i}(a+b^{T}(X_{i}-x))-C(a+b^{T}(X_{i}-x)).$
The penalized local logistic estimator spelled out in (2) is a special case of this more general construction, with $N=1$ and $C(t)=\log(1+e^{t})$. This log-likelihood is again a concave objective function, so under further regularity conditions on $C$ and after calculations analogous to those of Sections S.1.3 and S.1.4, our general theory in Theorem 4 and Proposition 5 will apply and yield, after straightforward calculations, that $(\sqrt{k}(\widehat{a}_{n}(x)-\ell_{C}(x)),\tau_{n,k}(x)\sqrt{k}(\widehat{b}_{n}(x)-\nabla\ell_{C}(x)))$ has a nontrivial weak limit, where $\ell_{C}(x)=C^{\prime-1}\circ m(x)=C^{\prime-1}\circ g(\beta^{T}x)$. Since $\nabla\ell_{C}(x)$ belongs to $\operatorname{span}(\beta)$, this estimation technique provides a way of recovering the dimension reduction subspace in a multi-class setting under a GLM specification.
The argument of course extends in a straightforward fashion to the situation where $Y$ and the corresponding parameter mapping are multivariate: for example, a multinomial model with $K\geq 2$ classes has conditional probability mass function
$(y_{1},\ldots,y_{K-1})\mapsto\pi_{1}(x)^{y_{1}}\cdots\pi_{K-1}(x)^{y_{K-1}}\left(1-\sum_{j=1}^{K-1}\pi_{j}(x)\right)^{1-\sum_{j=1}^{K-1}y_{j}}$
where $y_{1},\ldots,y_{K-1}\in\{0,1\}$ and $\sum_{j=1}^{K-1}y_{j}\in\{0,1\}$. This is readily put in exponential form, resulting in a probability mass function of the type $y\mapsto h(y)\exp(y^{T}\theta(x)-C(\theta(x)))$ with $y=(y_{1},\ldots,y_{K-1})$ and $\theta(x)=(\pi_{1}(x),\ldots\pi_{K-1}(x))$ being $(K-1)-$dimensional. In such a situation one would use the estimator
$(\widehat{a}_{n}(x),\widehat{b}_{n}(x))=\operatorname*{arg\,max}_{(a,b)\in\mathbb{R}^{K-1}\times\mathbb{R}^{(K-1)\times p}}\{\mathcal{L}_{n}(a,b)-\lambda\|b\|_{1}\}$
with, for any $(a,b)\in\mathbb{R}^{K-1}\times\mathbb{R}^{(K-1)\times p}$,
$\mathcal{L}_{n}(a,b)=\sum_{i\in N_{k}(x)}Y_{i}^{T}(a+b^{T}(X_{i}-x))-C(a+b^{T}(X_{i}-x)).$
This makes it possible to construct nearest-neighbor dimension reduction methodologies for multinomial responses in either the nominal or ordinal models; see Chapter 8 in Agresti (2013) for a concise introduction to such models in the classical GLM setting.
@@ 3.1.1 Estimation of the dimension reduction matrix
For a given choice of $m$ (the number of $b$’s used to estimate $M$), a given value of $\lambda$, the penalization parameter in the optimization problem, and $k$ the number of neighbors, the computation of the dimension reduction matrix is straightforward, see Algorithm 1 below (the optimization uses the R function glmnet from the package of the same name).
@@
To estimate the directions featured in (3), we use either the pure nearest-neighbor logistic log-likelihood without penalization or its penalized version. For the latter, we require an optimal choice of $\lambda$ which balances the regularization and goodness-of-fit. We propose to select the same parameter $\lambda$ for all $x\in\mathcal{X}_{m}$ in order to alleviate the computational burden. We therefore select $\lambda$ by 10-fold cross-validation at the average point $x=\overline{x}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$. In other words, we divide the data into 10 randomly selected subsets of equal size, each subset serving as a validation set while the remaining subsets are used for model fitting. The fitted model is assessed on the validation set using the misclassification error, that is, the proportion of observations whose label is not correctly predicted, which is the empirical counterpart of the misclassification risk $\mathcal{R}(g)=\mathbb{P}(g(X)\neq Y)$ for a given classifier $g$; in this cross-validation procedure, an observation is labeled as 1 if and only if its predicted probability of success by the nearest-neighbor logistic log-likelihood estimator $\widehat{\pi}_{n}(\overline{x})=\operatorname{expit}(\widehat{a}_{n}(\overline{x}))$ exceeds $1/2$. This evaluation of the quality of the fitted model is then done across a sequence of $\lambda$ values: more precisely, we use the R function cv.glmnet from the glmnet package with type.measure=class, which computes the average misclassification error across all validation sets, and the regularization parameter $\lambda$ selected is the one minimizing this error.
@@ 3.1.2 Selecting the dimension of the reduction subspace
Unlike some dimension reduction approaches such as that of the seminal paper by Li (1991), we estimate the dimension of the reduction subspace using cross-validation on the underlying prediction problem, and not from a test on the eigenvalues of the outer product of gradients $\widehat{M}$. We refer to Bura and Yang (2011) and Portier and Delyon (2014) for more details on the rank testing procedures used to estimate the dimension of the reduction subspace. Our procedure first divides the data into a training set $(X_{\mathrm{train}},Y_{\mathrm{train}})$ and a testing set $(X_{\mathrm{test}},Y_{\mathrm{test}})$. The matrix $\widehat{M}$ and its $p$ orthogonal eigenvectors $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{p}$, in decreasing order according to their eigenvalues, are estimated from the training set following the procedure described in Algorithm 1. For every $d\in\{1,\ldots,p\}$, the first $d$ eigenvectors are then gathered in a matrix $\widehat{\beta}_{(1:d)}\in\mathbb{R}^{p\times d}$, the sets of covariates $X_{\mathrm{train}}$ and $X_{\mathrm{test}}$ are projected onto the sets of lower-dimensional covariates $X_{\mathrm{train}}\widehat{\beta}_{(1:d)}$ and $X_{\mathrm{test}}\widehat{\beta}_{(1:d)}$, and a classifier is learned based on the training set $(X_{\mathrm{train}}\widehat{\beta}_{(1:d)},Y_{\mathrm{train}})$. We use here mainly the knn nearest-neighbor classifier, that is, at a given point $x$, the result of the majority vote among the nearest neighbors of $x$ within the space of projections of the covariates in the training set (with ties broken at random). For this classifier, the training step merely consists of storing the covariates in the training set along with their labels, which will form the basis for the vote at each point. In the real data analysis, we shall also compare our results with the Random Forest classifier, whose training step is nontrivial. The performance of the chosen classifier is then evaluated on the test set $(X_{\mathrm{test}}\widehat{\beta}_{(1:d)},Y_{\mathrm{test}})$, and the dimension retained is the one for which the classifier has the lowest misclassification risk.
This forward iterative procedure, summarized in Algorithm 2, strike a balance between dimension reduction and the preservation of relevant information for classification. We shall compare the results obtained with the situation where the correct dimension of the reduction subspace is known in order to assess the influence of the dimension selection step.
@@ 3.2 Simulation study
In the examples presented below, $(X_{i},Y_{i})_{i=1,\ldots,n}$ is a collection of independent and identically distributed random copies of the pair $(X,Y)$, and $X$ is a vector of independent centered and unit Gaussian covariates with dimension $p\geq 8$. We consider the following four examples:
Example 1. The response $Y$ has a Bernoulli distribution with parameter $\operatorname{expit}(X_{1}+\alpha=1)$.
Example 2. The response is $Y=\operatorname{sign}\{\sin(X_{1})+X_{2}^{2}+0.2\epsilon\}$, where $\epsilon\sim\mathcal{N}(0,1)$.
Example 3. The response is $Y=\operatorname{sign}\{(X_{1}+0.5)(X_{2}-0.5)^{2}+0.2\epsilon\}$, where $\epsilon\sim\mathcal{N}(0,1)$.
Example 4. The response is $Y=\operatorname{sign}\{\log(X_{1}^{2})(X_{2}^{2}+X_{3})+0.2\epsilon\}$, where $\epsilon\sim\mathcal{N}(0,1)$.
Example 1 is the simplest instance of a logistic regression model with a single relevant feature, while Examples 2, 3, and 4 are closely related to examples considered by Meng et al. (2020). We compare the proposed approach, with penalization (denoted by LLO$(\lambda>0)$) and without penalization (called LLO$(\lambda=0)$) with three existing competitors: SAVE and PHD, which are inverse regression techniques (Cook and Weisberg, 1991; Li, 1992) (both using the dr function from the R package of the same name) and POTD, relying on optimal transport (Meng et al., 2020). They are compared using, on the one hand, the distance between the estimated central subspace and the true central subspace using the Frobenius distance between the projection matrices on these subspaces, that is, $d(\mathcal{S}(\widehat{\beta}),\mathcal{S}({\beta}))=\|P_{\widehat{\beta}}-P_{\beta}\|_{F}$ with $P_{\beta}=\beta\beta^{T}$ for any orthogonal matrix $\beta$ of $p-$dimensional vectors. In each example, the central subspace is explicit in Example 1, the central subspace is spanned by the first vector of the canonical basis in $\mathbb{R}^{p}$, it is spanned by the first two vectors of this basis in Examples 2 and 3, and it is spanned by the first three vectors of this basis in Example 4. As such, Examples 1 to 4 are ranked in order of complexity of the dimension reduction problem.
(a) Distance to central subspace
(b) Misclassification risk
(c) Distance to central subspace
(d) Misclassification risk
Figure 1: Simulation study – Distance to the central subspace (left, on the log scale) and misclassification risk (right) in Example 1 (top) and Example 2 (bottom), averaged over $N=1000$ replications in each situation. The covariate has dimension $p=8$ and the dimension $d$ is chosen as the dimension of the correct population central subspace (i.e. $d=1$ in Example 1 and $d=2$ in Example 2). In the right-hand panels, the misclassification risk represented relates to the nearest neighbor majority vote classifier using the set of projected covariates on the estimated subspace produced by each method; in addition, “Full” denotes this classifier on the full, non-projected set of covariates, and “Oracle” denotes this classifier using the covariates projected on the correct population central subspace.
@@
(a) Distance to central subspace
(b) Misclassification risk
(c) Distance to central subspace
(d) Misclassification risk
Figure 2: Simulation study – Distance to the central subspace (left) and misclassification risk (right) in Example 3 (top) and Example 4 (bottom), averaged over $N=1000$ replications in each situation. The covariate has dimension $p=8$ and the dimension $d$ is chosen as the dimension of the correct population central subspace (i.e. $d=2$ in Example 3 and $d=3$ in Example 4). In the right-hand panels, the misclassification risk represented relates to the nearest neighbor majority vote classifier using the set of projected covariates on the estimated subspace produced by each method; in addition, “Full” denotes this classifier on the full, non-projected set of covariates, and “Oracle” denotes this classifier using the covariates projected on the correct population central subspace.
@@
(a) Distance to central subspace
(b) Misclassification risk
Figure 3: Simulation study – Distance to the central subspace (left) and misclassification risk (right) in Example 4, averaged over $N=1000$ replications of a sample of size $n=1000$, as a function of the dimension $d\in\{1,\ldots,6\}$ of the estimated central subspace and $p\in\{8,16,32,64\}$ of the full covariate space. In the right-hand panels, the red dashed line corresponds to the nearest-neighbor classifier with $d=p$, and the blue dashed line corresponds to this classifier using the covariates projected on the correct population central subspace.